3. [BC#1.5.8] Show that for any complex numbers zi,22 E C, we have = a1 +bi, etc.), or you can appeal to po- This can be done directly ( lar/exponential form(s)...
How to prove this
Answer #1

To show:

\begin{vmatrix} Z_{1}Z_{2} \end{vmatrix}=\begin{vmatrix} Z_{1} \end{vmatrix}\times \begin{vmatrix} Z_{2} \end{vmatrix}

Proof:

Let

Z_{1}=a_{1}+b_{1}i

Z_{2}=a_{2}+b_{2}i

So,

Z_{1}Z_{2}=(a_{1}+b_{1}i)\times (a_{2}+b_{2}i)

=(a_{_{_{1}}}a_{2}-b_{1}b_{2})+(a_{1}b_{2}+a_{2}b_{_{_{1}}})i

So,

\begin{vmatrix} Z_{1}Z_{2} \end{vmatrix}=\sqrt{(a1a2-b1b2)^{2}+(a1b2-a2b1)^{2}}

=\sqrt{a1^{2}a2^{^{2}}+b1^{2}b2^{2}-2a1a2b1b2+a1^{2}b2^{2}+a2^{2}b1^{2}+2a1b2a2b1}

=\sqrt{a1^{2}(a2^{2}+b2^{2})+b1^{2}(a2^{2}+b2^{2})}

=\sqrt{(a1^{2}+b1^{2})(a2^{2}+b2^{2})}

=\sqrt{(a1^{2}+b1^{2})}\sqrt{(a2^{2}+b2^{2})}}

=\begin{vmatrix} Z_{1} \end{vmatrix}\times \begin{vmatrix} Z_{_{_{2}}} \end{vmatrix}

This completes the proof.